OpenStudy (anonymous):

x squared +8x+20=0 help please!!

OpenStudy (anonymous):

what year level are you?

OpenStudy (anonymous):

x^2 +8x +20

OpenStudy (anonymous):

this has a complex soln

OpenStudy (anonymous):

factor it. (x+10)(x-2) = o x = -10 and 2

OpenStudy (anonymous):

grade 10

OpenStudy (anonymous):

ummmm ... @iamtheos its not x^2 +8x-20=0

OpenStudy (anonymous):

sorry read it wrong. have you used the quadratic equation?

OpenStudy (anonymous):

yes i did

OpenStudy (anonymous):

This has complex roots ... if you are only grade ten then this is insolvable for you as u will have a negative determinate (sqrroot) in your quadratic eqn

OpenStudy (anonymous):

well i just entered 10 grade

OpenStudy (anonymous):

x = -4-2 i x = -4+2 i lol u could blow ur teacher away by telling them that its complex roots are these lol, this is a 4 unit topic